OpenStudy (anonymous):

What is the ratio of x to y? 3x = 8y

ganeshie8 (ganeshie8):

divide "y" both sides

ganeshie8 (ganeshie8):

you would get : 3x/y = 8

ganeshie8 (ganeshie8):

next divide 3 both sides

OpenStudy (anonymous):

How do i get 3x/y=8

ganeshie8 (ganeshie8):

3x = 8y divide "y" both sides 3x = 8y /y /y

ganeshie8 (ganeshie8):

"y" cancels on right side, eh ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

left hand side, it just stay in the denominator : 3x/y = 8

ganeshie8 (ganeshie8):

fine, so far ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

next divide 3 both sides, and you're done !

ganeshie8 (ganeshie8):

3x/y = 8 /3 /3 x/y = 8/3

OpenStudy (anonymous):

Thank you.

ganeshie8 (ganeshie8):

u wlc :)